Suppose the short run productiin function can be represented by Q=60,000L2-1000L3. then, determine

A) the level labor employement that maximizes the level of output

B) the level of employement that maximizes APL and the maximum APL


1
Expert's answer
2022-01-10T14:43:11-0500

"Q=60000L^2-1000L^3,"

"Q=0,"

"L=\frac{60000}{1000}=60,"


"APL=Q'=120000L-3000L^2,"

"APL=0,"

"L=\frac{120000}{3000}=40,\implies"

"Q=60000\cdot 40^2-1000\cdot 40^3=32000000."